# SPECIAL RELATIVITY

### v9

@@ 4.2 The ”Plane Wave” Emitter
This problem in special relativity can be adequately treated by an approach which uses the absolute time coordinatization. We will demonstrate that when one has a transversely moving mirror with finite aperture and a plane wave of light is falling normally onto the mirror, there is a deviation of the energy transport for the reflected light. In this case, the effect of aberration of light results directly from the time-dependence of the position of the moving mirror, since a finite aperture has to be considered. What must be recognized is that in the time-dependent emitter-mirror problem, the solution involves light beams with different frequencies. It comes out quite natural that the result will depend on the velocity vector: in other words, we account to the influence of the Doppler effect. For a transversely moving mirror with finite aperture we cannot neglect the angular frequency dispersion, which is an effect of the first order in $v/c$.
@@
We already pointed that Chapter 3 is probably the most important chapter in this book. To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature. Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain metric Eq. 11. This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements.
@@
There is an intuitively plausible way to understanding why authors of textbooks conclude that there is no aberration for light reflected from a transversely moving mirror. A comparison between the conventional aberration of light theory and the sound theory might help here. Bellow in the Section 4.15 we shall consider the sound emitter at rest in the atmosphere frame. We shall consider the case of a mirror, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for reflected beam. The group velocity is equal to the phase velocity. A plane wave is falling normally on the mirror, and generates an oblique sound beam. The (diagonal) wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest. According to textbooks, there are important parallels between aberration of sound and light in the first order approximation. This is misconception. There is a fundamental difference between velocity of light and velocity of sound. The satisfactory treatment of the aberration of light phenomena should be based on the pseudo-Euclidean space-time geometry.
@@ 4.6 Way to Solve the Emitter-Mirror Problem in the (3+1) Space and Time
The essence of Abbe’s approach, in our case of interest, is that one regards the mirror as a kind of a diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams constituted by plane waves. Each of these beams corresponds to one of the Fourier components into which the reflected power of the mirror can be resolved. The finite-aperture mirror is a non-periodic object. It gives an infinite number of diffracted beams forming a continuum.
@@
The $\vec{k}$ vectors shown in Fig.4 represent the propagation vector of the incident plane wave $\vec{k}_{i}$, which is assumed to be directed perpendicularly to the surface. The vectors $\vec{k_{s}}^{(+)}$ and $\vec{k_{s}}^{(-)}$ are added to indicate the scattered light. The Bragg condition $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ shows how the direction of the incident and scattered wave are related. The first-order maxima dominate due to the fact that light is being scattered from a sinusoidal grating, rather than a set of discrete planes (grooves).
@@
When the scatterer wave is a progressive wave rather than a fixed modulation, the frequency of the scattered wave is different from that of the incident wave. This fact is interpreted as a Doppler effect, since the reflection is from a moving, rather than a stationary, set of waves. In the case of a transversely moving grating, light in the diffraction maxima undergoes a Doppler shift resulting from the fact that it has been reflected from moving waves with wavenumber vectors $\vec{K}_{\perp}$.
Let us demonstrate that the new terms that have to put into the field equations due to use of absolute time coordinatization lead to the prediction of the Doppler effect. We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c+k_{z}^{2}-\omega^{2}/c^{2}=0$. The wavenumber vector of the radiated plane wave is fixed by initial conditions. In fact, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance. From this dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$.
@@
The following important detail of such ”single inertial frame” description can hardly be emphasize enough. If the source of light is at rest and the mirror is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the absolute time coordinatization should be applied and kept in a consistent way for both incoming and scattered waves (Fig. 4). In our previous discussion of absolute time coordinatization we learned that the emitter at rest must be in the same time described by Maxwell’s electrodynamics. A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{i}^{2}-\omega_{i}^{2}=0$. From the initial conditions we find that $\vec{k}_{i}=\vec{e}_{z}k_{z}$, $\omega_{i}=ck_{z}$. The contradiction, however disappears if we perform geometrical analysis of light reflection. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$-axis the dispersion equation in the absolute time coordinatization will have the same (diagonal) form $k_{z}^{2}-\omega^{2}=0$ for the incident wave. We worked out a very simple case in order to understand all the physical principle very clearly. In Section 5.11 we shall analyze the aberration of light with incoming wave which has rate of change of phase with respect to $x$-axis.
@@
Figure 5: Transversely moving mirror with a small aperture at normal incidence. When a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for the reflected light beam. This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of the light waves reflected from the mirror. As a result, the velocity of the energy transport is not equal to the phase velocity.
One of the most important conclusions of the foregoing discussion is a remarkable prediction on the theory of the aberration of light, concerning the deviation of the energy transport for light reflected from a mirror moving transversely. Namely, when a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for reflected light beam (see Fig. 5). This phenomenon can be regarded as a simple consequence of the Doppler effect.
The argument that in the process of reflection from a transversely moving mirror the direction of propagation is not given by the normal to the wavefront is considered erroneous in literature (4). This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. The direction of the energy transport has an exact objective meaning i.e. is convention-invariant. However, the phase front orientation (i.e. the plane of simultaneity in the judgement of an observer) has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this orientation could be ascertained.
Since the phase front orientation does not exist as physical reality within the angular range $v/c$, a question arises: why do we need to account for the exact phase front orientation in our electrodynamics calculations? The answer is that when the evolution of the radiation beam is treated according to the single inertial frame of reference, one will experience that phase front orientation remains unvaried: this has no objective meaning but is used in the analysis of the electrodynamics problem. A comparison with a gauge transformation in Maxwell’s electrodynamics might help here. Even if the phenomena are quite different, the common mathematical formulation permits us to draw this analogue.
@@ 4.7 Discussion
Physicists who try to understand the situation related to the use of the textbook approach in the aberration of light phenomena, are often troubled by the fact that in the situation where there is a mirror moving normally to its surface the reasoning presented in textbooks is correct. The reflection from the mirror is analyzed in two Lorentz reference frames. It is interesting to note that, for this case, relativistic kinematics correctly predicts the light frequency variation on reflection from a moving mirror. At this point a reasonable question arises: why the same method gives incorrect result in the case of reflection of light from a mirror moving tangentially to its surface?
It is not difficult to see that the peculiarity of the situation where there is an infinite mirror moving normally to its surface is that here the ”mirror motion” is a real observable effect. Indeed, even if the mirror is infinite the problem is time dependent. In this situation, a velocity of the infinite mirror has physical meaning. Thus, for normal motion, the infinite mirror may be a useful concept for analysis of the Doppler effect. This method was incorrectly extended to the case of tangential motion.
The second peculiarity of the collinear geometry where there is a mirror moving normally along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization. We shall discuss this subject further in the Chapter 10.
@@ 4.8 Large Aperture Mirror
In this chapter our discussion is limited to the region of problem parameters, in which we forget about the emitter edges. Although this aberration of light theory is just an approximation, it is a very great importance practically. We shall also discuss the situation where the transverse size of the emitter is very small relative to the transverse size of the moving mirror. It should be note that this situation is not realized in the stellar aberration measurements. We shall work out this case in order to understand all the physical principles very clearly. It is easy to show that the deviation of the energy transport is absent in this case, Fig. 6. The direct approach to moving mirror problem uses the Fourier transform methods. We now consider another method of calculating the aberration of light effect - we want to illustrate the great variety of possibilities. The way of thinking that made the law about the behavior of light reflected from a large aperture mirror evident is based on Babinet’s principle. It is well known that, when light comes through a hole of a given shape, made in an opaque screen, the distribution of intensity after the hole (i.e. the diffraction pattern) is the same as in the case when the hole is replaced by sources (dipoles) uniformly distributed over the hole. In other words, the diffracted plane wave from a hole, or from a source with the same shape of the hole are the same.
Figure 6: Transversely moving mirror with a large aperture at normal incident. When a beam of light is falling normally on the mirror, there is no deviation of the energy transport for the reflected light beam. The velocity of the energy transport is equal to the phase velocity.
This is a particular case of Babinet’s principle, which states that the sum of diffraction fields behind two complementary opaque screens is the incident wave. We know from this principle, that the solution we have found using Abbe’s approach also corresponds to that for large aperture hole in a moving opaque screen. We see clearly that there is no electromagnetic interaction of a light beam with screen. Indeed light beam is not scattered by the hole edges. Does this discussion about large aperture hole have any meaning? To see whether it does, we should remember about the Babinet’s principle. Here we only wish to show how easy the law of reflection from a large aperture mirror can be found with the help of the Babinet’s principle.
@@ 4.9 Analysis of Transmission through a Hole in a Opaque Screen
Above we demonstrated that when one has a small aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror. The problem to be considered in this section is of more practical importance. We now consider the case of a screen, in the lab frame, moving with velocity $v$ along its surface. It is generally believed that there is no deviation of the energy transport for light transmitted through a hole in the moving opaque screen, Fig. 7.
However, there is a common mistake made in relativistic optics, connected with aberration effects from a transversely moving screen containing a hole. We describe the system using, again, a Fourier transform method similar to that considered above. The screen containing a hole is a kind of diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which a transmitted light beam can be resolved.
@@
Figure 7: Aberration of light in an inertial frame of reference. Transversely moving screen which has a hole in it. According to textbooks, a monochromatic plane wave of light is falling normally on the screen and generates a transmitted oblique beam. There is no deviation of the energy transport for the transmitted oblique light beam. The velocity of the energy transport is equal to the phase velocity.
@@
Figure 9: Aberration of light in an inertial frame of reference. Transversely moving screen which has a hole in it. A monochromatic plane wave of light is falling normally on the screen and generates a transmitted light beam. The Doppler effect is responsible for the angular frequency dispersion of the light waves transmitted through the hole. As a result, the velocity of the energy transport is not equal to the phase velocity.
According to our approach, there is a remarkable prediction of the theory of aberration of light concerning the deviation of the energy transport for light transmitted through a hole in a moving screen. Namely, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through the hole (see Fig. 9).

### v10

@@ 4.1 The ”Plane Wave” Emitter
This problem in special relativity can be adequately treated by an approach that uses absolute time coordinatization. We will demonstrate that when one has a tangentially moving mirror with a finite aperture and a plane wave of light is falling normally onto the mirror, there is a deviation of the energy transport for the reflected light. In this case, the effect of aberration of light results directly from the time-dependence of the position of the moving mirror, since a finite aperture has to be considered. What must be recognized is that in the absolute time coordinatization, the wave equation will depend on the velocity vector. It comes out quite naturally that the solution involves light beams with different frequencies. In other words, we account for the influence of the Doppler effect. For a tangentially moving mirror with finite aperture, we cannot neglect the angular frequency dispersion, which is an effect of the first order in $v/c$.
@@
We already pointed out that Section 3 is probably the most important section in this article. To our knowledge, neither the operational interpretation of absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature. Using the Galilean boost: we substitute $x\to x-vt$ while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain metric Eq. 11. This metric describes the electrodynamics of the moving mirror from the viewpoint of the inertial observer measurements.
@@
There is an intuitively plausible way to understand why authors of textbooks conclude that there is no aberration for light reflected from a tangentially moving mirror. A comparison between the conventional aberration of light theory and the sound theory might help here. In Section 4.13 we shall consider the sound emitter at rest in the atmosphere frame. We shall consider the case of a mirror, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for the reflected beam. The group velocity is equal to the phase velocity. A plane wave falls normally on the mirror and generates an oblique sound beam. The (diagonal) wave equation of the moving mirror in the atmosphere frame is the same as the wave equation of the emitter at rest. According to textbooks, there are important parallels between the aberration of sound and light in the first-order approximation  [^18] We should make one further remark about two errors in solving problems concerning reflection from a mirror moving tangentially to its surface. These errors are not independent. According to textbooks, the metric of the moving mirror is diagonal as the metric of the ”plane wave” emitter at rest, and the group velocity is equal to the phase velocity. So independently of the mirror size, there is no deviation of the energy transport for the reflected light beam.. This is a misconception. There is a fundamental difference between the velocity of light and the velocity of sound. The satisfactory treatment of the aberration of light phenomena should be based on the pseudo-Euclidean space-time geometry.
@@ 4.5 Way to Solve the Emitter-Mirror Problem in the (3+1) Space and Time
The essence of Abbe’s approach, in our case of interest, is that one regards the mirror as a kind of diffraction grating that breaks up the incident beam of the plane wave into a number of diffracted beams constituted by plane waves. Each of these beams corresponds to one of the Fourier components into which the reflected power of the mirror can be resolved. The finite-aperture mirror is a non-periodic object. It gives an infinite number of diffracted beams forming a continuum.
@@
The $\vec{k}$ vectors shown in Fig.4 represent the propagation vector of the incident plane wave $\vec{k}_{i}$, which is assumed to be directed perpendicularly to the surface. The vectors $\vec{k_{s}}^{(+)}$ and $\vec{k_{s}}^{(-)}$ are added to indicate the scattered light. The Bragg condition $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ shows how the direction of the incident and scattered wave are related. The first-order maxima dominates because light is being scattered from a sinusoidal grating, rather than a set of discrete planes (grooves).
@@
When the scatterer wave is a progressive wave rather than a fixed modulation, the frequency of the scattered wave is different from that of the incident wave. This fact is interpreted as a Doppler effect, since the reflection is from a moving, rather than a stationary, set of waves. In the case of a tangentially moving grating, light in the diffraction maxima undergoes a Doppler shift resulting from the fact that it has been reflected from moving waves with wavenumber vectors $\vec{K}_{\perp}$.
Let us demonstrate that the new terms that have to be put into the field equations due to the use of absolute time coordinatization lead to the prediction of the Doppler effect. We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c+k_{z}^{2}-\omega^{2}/c^{2}=0$. The wavenumber vector of the radiated plane wave is fixed by initial conditions. In fact, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance. From this dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$  [^19] We shall now give one more derivation of this same interesting and important result. We note the very close correspondence between angular and frequency response of moving diffraction gratings and the Raman scattering [25]. Since the theory of Bragg diffraction applies to light scattering by sound waves in liquids and solids, it is not surprising that we can obtain our results from the quantum theory of light scattering by phonons. This scattering process is also known as Brillouin scattering. The quantum theory leads to equation $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ by the requirement that the momentum is conserved between interacting photon and phonon. The momentum balance for scattering involving the emission or absorption of phonon leads to the Bragg condition. The frequency shift that occurs for light scattering from sound waves comes about as a result of the Doppler effect. Quantum mechanically, it is a consequence of the conservation of energy between the participating particles. In contrast, authors of special relativity textbooks conclude that there is no Doppler effect for light reflected from a tangentially moving grating. They discuss about a tangentially moving mirror but it is also true for grating. The difficulty we speak of is associated with the generally accepted assumption that the metric of the moving grating (mirror) is the same (i.e. Minkowski metric) as the metric of the emitter at rest. .
@@
The following important detail of such a ”single inertial frame” description can hardly be emphasized enough. If the source of light is at rest and the mirror is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the absolute time coordinatization should be applied and kept in a consistent way for both incoming and scattered waves (Fig. 4). In our previous discussion of absolute time coordinatization, we learned that the emitter at rest must be at the same time described by Maxwell’s electrodynamics. A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{i}^{2}-\omega_{i}^{2}=0$. From the initial conditions we find that $\vec{k}_{i}=\vec{e}_{z}k_{z}$, $\omega_{i}=ck_{z}$. The contradiction, however, disappears if we perform a geometrical analysis of light reflection. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$-axis the dispersion equation in the absolute time coordinatization will have the same (diagonal) form $k_{z}^{2}-\omega^{2}=0$ for the incident wave. We worked out a very simple case to understand all the physical principles very clearly. In Section 5.10 we shall analyze the aberration of light with an incoming wave which has a rate of change of phase with respect to the $x$-axis.
@@
Figure 5: Tangentially moving mirror with a small aperture at normal incidence. When a plane wave of light falls normally on the mirror, there is a deviation of the energy transport for the reflected light beam. The explanation of the effect of aberration is based on the use of absolute time coordinatization.
One of the most important conclusions of the foregoing discussion is a remarkable prediction on the theory of the aberration of light, concerning the deviation of the energy transport for light reflected from a mirror moving tangentially to its surface. Namely, when a plane wave of light (in the absolute time coordinatization) falls normally on the moving mirror, there is a deviation of the energy transport for the reflected light beam (see Fig. 5). This phenomenon can be regarded as a simple consequence of the Doppler effect [^20] Doppler effect is just a mathematical description in an intermediate step: when a complete calculation is made, only the geometrical (aberration) effect takes place..
The argument is that in the process of reflection from a tangentially moving mirror the direction of propagation is not given by the normal to the wavefront is considered erroneous in literature  [^21] There is a common misconception that the radiation wavefront orientation has objective meaning. To quote Norton [26]: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame then that must be true in all inertial frame.”. This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. The direction of the energy transport has an exact objective meaning i.e. is convention-invariant. However, the phase front orientation (i.e. the plane of simultaneity in the judgment of an observer) has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this orientation could be ascertained.
Since the phase front orientation does not exist as physical reality within the angular range $v/c$, a question arises: why do we need to account for the exact phase front orientation in our electrodynamics calculations? The answer is that when the evolution of the radiation beam is treated according to the single inertial frame of reference, one will experience that phase front orientation remains unvaried: this has no objective meaning but is used in the analysis of the electrodynamics problem. A comparison with a gauge transformation in Maxwell’s electrodynamics might help here.
@@ 4.6 Discussion
Physicists who try to understand the situation related to the use of the textbook approach in the aberration of light phenomena are often troubled by the fact that in the situation where there is an infinite mirror moving normally to its surface the reasoning presented in textbooks is correct. The reflection from the mirror is analyzed in two Lorentz reference frames. In this case, relativistic kinematics correctly predicts the light frequency variation on reflection from a moving mirror. At this point, a reasonable question arises: why does the same method give incorrect results in the case of the reflection of light from a mirror moving tangentially to its surface?
It is not difficult to see that the peculiarity of the normally moving (mirror) source is that here the velocity is perpendicular to the plane of oscillating elementary sources (dipoles). Thus, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization. In the case of collinear geometry, a motion of the mirror, according to the theory of relativity, influences the relativistic kinematics terms of the second order only. Let us now consider a frequency measurement in the context of special relativity. Suppose that we have a Fabry-Perot interferometer (or a grating spectrometer). The frequency measured in this case is (standing wave) wavelength. The same statement is also true for all cases of light frequency measurements. Frequency measurements automatically obey the relationship between light frequency and interference. In contrast, the deviation of the energy transport direction is a geometrical effect. There is a fact governing all interference phenomena. It states that the interference pattern is independent of the chosen metric. In other words, according to the space-time geometry, phase is 4D invariant. For collinear geometry, Minkowski metric and metric Eq. 11 give the same prediction (see Chapter 13 for more details).
The next thing that we have to discuss is the mirror size. The second peculiarity of the situation where there is an infinite mirror moving normally to its surface is that here the ”mirror motion” is a real observable effect. Indeed, even if the mirror is infinite the problem is time dependent. In this situation, the velocity of the infinite mirror has physical meaning. Thus, for normal motion, the infinite mirror may be a useful concept for the analysis of the Doppler effect.
@@ 4.7 Large Aperture Mirror
In this section, our discussion is limited to the region of problem parameters, in which we forget about the emitter edges. Although this aberration of light theory is just an approximation, it is of very great importance practically. We shall also discuss the situation where the transverse size of the emitter is very small relative to the transverse size of the moving mirror. It should be noted that this situation is not realized in the stellar aberration measurements. We shall work out this case to understand all the physical principles very clearly. It is easy to show that the deviation of the energy transport is absent in this case, Fig. 6. The direct approach to the moving mirror problem uses the Fourier transform method.
Figure 6: Tangentially moving mirror with a large aperture at normal incident. When a beam of light falls normally on the mirror, there is no deviation of the energy transport for the reflected light beam.
@@ 4.8 Analysis of Transmission through a Hole in a Opaque Screen
Above we demonstrated that when one has a finite aperture mirror moving tangentially to its surface and the plane wave of light falls normally on the mirror, there is an aberration (deviation of the energy transport) for light reflected from the mirror. The problem to be considered in this section is of more practical importance. We now consider the case of a screen, in the lab frame, moving with velocity $v$ along its surface. It is generally believed that there is no deviation of the energy transport for light transmitted through a hole in the moving opaque screen, Fig. 7.
However, there is a common mistake made in relativistic optics, connected with aberration effects from a tangentially moving screen containing a hole. We describe the system using, again, a Fourier transform method similar to that considered above. The screen containing a hole is a kind of diffraction grating that breaks up the incident beam of the plane wave into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which a transmitted light beam can be resolved.
@@
Figure 7: Aberration of light in an inertial frame of reference. Tangentially moving screen which has a hole in it. According to textbooks, a monochromatic plane wave of light falls normally on the screen and generates a transmitted oblique beam.
@@
Figure 9: Aberration of light in an inertial frame of reference. Tangentially moving screen which has a hole in it. A monochromatic plane wave of light falls normally on the screen and generates a transmitted light beam. The explanation of the effect of aberration is based on the use of absolute time coordinatization.
According to our approach, there is a remarkable prediction of the theory of aberration of light concerning the deviation of the energy transport for light transmitted through a hole in a moving screen. Namely, when one has a tangentially moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through the hole (see Fig. 9).